Sum of the ages of fahad and ahmad is 38.Seven years ago, fahad was three times as old as ahmad.find their present ages
Answer by Abbey(339)

Fahad = x
Ahmad=y
Fahad seven years ago = x-7
Ahmad seven years ago = y-7
x+y=38 OR y=38-x
Seven years ago:
(x-7)=3(y-7)
x-7=3y-21
x+14=3y
Substitute 38-x for y
x+14=3(38-x)
x+14=114-3x
add 3x-14 to both sides
4x=100
x=25
so y=13
This is correct because 7 years ago Fahad was 18 and ahmad was 6